OpenStudy (anonymous):

find the maximum and minimum value of y=3x^3 - 4x +2

OpenStudy (anonymous):

\[\frac{ d (3x^3 - 4x +2)}{ dx} = 9x^2-4\]

OpenStudy (anonymous):

9x^2-4=0 x= 2/3 x=-2/3 nw wat to do

OpenStudy (amistre64):

as in the range of the function? its a cubic; odd integer degrees rise towards infinity and dive towards negative infinity

OpenStudy (amistre64):

if your trying to determine the bumps and dips; take the derivative a second time, insert the zeros from that, and see if the signs are positive or negative

OpenStudy (amistre64):

positive opens up; U which translates as a min negative opens Down n, which translates as a max